OpenStudy (anonymous):

f(x)=3x2+5x find f(-2)

OpenStudy (promisespromises):

Just plug -2 in for all x in the equation.

OpenStudy (anonymous):

Its mean x=-2. change x by -2. 3(-2)2+5(-2)

OpenStudy (unklerhaukus):

\[f(x)=3(x)^2+5(x) \] \[f(-2)=3(-2)^2+5(-2)\\\quad\qquad= \]